OpenStudy (anonymous):

Find the surface area of a cylinder with diameter 22 in and height 15 in.

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (anonymous):

surface area=2*pi*r(r+h) since diameter =22 ,so r=11 putting values in formulae of surface area we get 12584/7=1796.99